# WAVE: Weighted Autoregressive Varying Gate for Time Series Forecasting

### v3

@@ 1 Introduction
In recent years, autoregressive (AR) decoder-only Transformer-based models (Vaswani, 2017; Radford, 2018) have been widely used in sequence modeling tasks across fields such as NLP (Brown et al., 2020; Touvron et al., 2023), CV (Chen et al., 2020; Esser et al., 2021; Chang et al., 2022), and audio (Borsos et al., 2023). This structure is well-suited for various sequential generation and prediction tasks. However, in typical sequence modeling tasks like time series forecasting (TSF), there has been less exploration of this architecture compared to other structures. Most of the best-performing recent TSF models are encoder-only Transformers (Liu et al., 2024a; Nie et al., 2022), MLPs (Das et al., 2023; Lu et al., 2024), or even linear models (Zeng et al., 2023; Xu et al., 2024). The few relevant discussions mainly focus on using pretrained autoregressive LLMs or similar structures for few-shot and zero-shot prediction (Gruver et al., 2023; Jin et al., 2024; Das et al., 2024; Liu et al., 2024b), with little research directly evaluating their TSF performance in end-to-end training. Therefore, this paper will first briefly demonstrate that with appropriate tokenization and training methods, a basic AR Transformer is enough to achieve results comparable to the state-of-the-art (SOTA) baselines, as shown in Fig. 1.
@@ 2.2 Appropriate tokenization for autoregressive forecasting
Recently, most time series forecasting research utilizes encoder-decoder or encoder-only Transformers for TSF (Li et al., 2019b; Zhou et al., 2021; Wu et al., 2021; Nie et al., 2022; Liu et al., 2024a), with limited focus on end-to-end decoder-only autoregressive Transformer because of error accumulation issue. For long-term forecasts, The autoregressive Transformers requires iteratively doing one-step prediction, leading to error accumulation and higher MSE compared to non-autoregressive models that generate the entire forecast at once.
@@ 2.5 Inspiration: decoupling the short-term impact
However, applying exponential decay to the gated linear attention AR weights may not align with the needs of TSF, which often involves stable periodic patterns alongside short-term impacts. TSF data frequently exhibit seasonal effects that differ from the transient long-term effects in NLP. These seasonal effects are stable and persist across the temporal dimension without decaying. As shown in Figure LABEL:fig:exp_decay_strategies, in scenarios involving seasonal effects, sudden changes, and local effects, relying solely on exponential decay in gated linear attention is not the most suitable approach for modeling TSF. Decoupling local effects into short-term MA weights, allowing the AR term to focus on modeling the seasonal and long-term effects it handles best, would be a better solution for TSF.
@@ 2.7 Indirect MA weight generation
With $\mathbf{\Theta}=\mathbf{B}\cdot(\mathbf{I}-\mathbf{B})^{-1}$, as long as the indirectly generated $\mathbf{\Theta}$ accurately reflects the characteristics of the MA weights we want, we can use $\sum_{j=1}^{t-1}\bm{\beta}_{t-1,j}\odot\bm{r}_{j}$ as $\bm{o}^{\text{MA}}_{t}$. Since $\bm{r}_{j}$ is known after computing $\bm{o}^{\text{AR}}_{t}$, linear attention can be used to compute $\bm{o}^{\text{MA}}_{t}$ without increasing the time complexity. To ensure the implicitly generated $\mathbf{\Theta}$ from $\mathbf{B}$ captures the desired MA properties, we must carefully design how $\mathbf{B}$ is generated. The invertibility of $(\mathbf{I}-\mathbf{B})^{-1}$ is guaranteed since $\mathbf{B}$ is strictly lower triangular. To efficiently compute the generated weights, we use the $\bm{\beta}_{t-1,j}=\phi_{q}^{\text{MA}}(\bm{q}^{\text{MA}}_{t-1})\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})$ to generate $\mathbf{B}$, similar to linear attention. Previous dynamic ARMA models in statistics often update MA weights based on observations (Grenier, 1983; Azrak & Mélard, 2006), so we derive $\bm{q}^{\text{MA}}_{t-1}$ and $\bm{k}^{\text{MA}}_{j}$ by multiplying the attention input $\bm{x}_{\cdot}$ with $\mathbf{W}^{\text{MA}}_{q}$ and $\mathbf{W}^{\text{MA}}_{k}$. Now, the effectiveness of MA weights lies in selecting the most suitable functions $\phi_{q}^{\text{MA}}(\cdot)$ and $\phi_{k}^{\text{MA}}(\cdot)$.
@@ 2.8 Selection of $\phi(\cdot)$ and characteristics of implicit MA weights
The MA term models short-term effects and local temporal relationships, so we want the implicit $\mathbf{\Theta}$ to follow a pattern where elements near the diagonal have larger absolute values, and those farther away gradually decrease. The expanded form of $\mathbf{\Theta}$ is given by $\mathbf{\Theta}=\mathbf{B}\cdot(\mathbf{I}-\mathbf{B})^{-1}=\mathbf{B}+\mathbf{B}^{2}+\mathbf{B}^{3}+\cdots$. The elements along the diagonal direction in $\mathbf{B}$ continually accumulate as products into the elements below them in $\mathbf{\Theta}$. Since $\mathbf{B}$ is strictly lower triangular, the elements of the subdiagonal in $\mathbf{\Theta}$ remain constant, while the elements further down progressively accumulate additional terms formed by the product of different $\beta_{\cdot}$ elements above. Assuming $\beta_{\cdot}$ follows a distribution and simplifying by setting each $\beta_{\cdot}$ to the distribution mean $b$, the elements of $\mathbf{\Theta}$ can be expressed as:
@@
We tested various activation function combinations for $\phi_{q}^{\text{MA}}(\cdot)$ and $\phi_{k}^{\text{MA}}(\cdot)$ to generate $\bm{\beta}_{t-1,j}=\phi_{q}^{\text{MA}}(\bm{q}^{\text{MA}}_{t-1})\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})$ values, as shown in Fig. LABEL:fig:vis_phi_selection. We used the sigmoid function $\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})=\sigma(\alpha\bm{k}^{\text{MA}}_{j}/\sqrt{d})$ to obtain values between 0 and 1, where $\alpha=0.05$  [^2] In the key activation, $\alpha$ controls the variance of each row in the $\mathbf{B}$ matrix, indirectly influencing the amount of long-term information (lower left) in the MA weights $\mathbf{\Theta}$. Increasing $\alpha$ would make the MA weights focus more on modeling long-term information. However, since we want the AR weights to handle the long-term component, we set $\alpha$ to a relatively small value. This explains why the rows of the $\mathbf{B}$ matrix appear smooth in the visualization. Refer to Fig. LABEL:fig:vis_phi_selection_alpha_selection for more details on $\alpha$, and see Fig. LABEL:fig:vis_phi_selection_positive for the effects of reversed positive $\phi_{q}$. and $\sqrt{d}$ are scaling factors to maintain small absolute values. Then, we selected a function $\phi_{q}^{\text{MA}}(\cdot)$ to make the product negative. We ultimately chose $\phi^{\text{MA}}_{q}(\bm{q}^{\text{MA}}_{t})=-\text{LeakyReLU}(-\bm{q}^{\text{MA}}_{t}/\sqrt{d})$ with a negative slope of 0.02. The inner negative sign maintains directional consistency (for later parameter sharing), and the outer negative sign encourages a negative output.
@@
Fig. LABEL:fig:vis_phi_selection shows that LeakyReLU provides a balanced lag weight pattern. Unlike ReLU and Sigmoid, which only output values of the same sign, LeakyReLU offers some relaxation while keeping most values negative. This adds flexibility by enabling the desired negative smoothing effect of the MA term, with occasional positive values to enhance modeling flexibility.
@@ 3 Experiments
We conducted comprehensive experiments on 12 widely-used TSF datasets, including Weather, Solar, Electricity (ECL), ETTs, Traffic, and PEMS. See §B.1 for detailed description of datasets.
@@
Baselines   We built AR Transformers using the five attention mechanisms from Table 1 and added MA terms to create WAVE attention for comparison in TSF tasks. Additionally, we included five recent SOTA baselines: FITS (Xu et al., 2024), iTransformer (Liu et al., 2024a), CATS (Lu et al., 2024), PatchTST (Nie et al., 2022), and DLinear (Zeng et al., 2023). We also used a simple channel-dependent encoder-only Transformer, modified by repeating the last input value (like NLinear) to address distribution shift. This model already surpasses older architectures like Autoformer (Wu et al., 2021) and Informer (Zhou et al., 2021), so we excluded these from our comparison.
In the main experiments, both pure AR and WAVE Transformers use a consistent setup: $m=3$ Transformer layers, 8 heads, and model dimension determined by a empirical method $d=16\sqrt{C}$, where $C$ is the number of series. We evaluate their performance using one-step prediction for each test datapoint, aligned with the baselines. Baseline hyperparameters are set to the reported values from their original papers. For more details on hyperparameters and implementation, see §B.2.
We ran all models on all datasets for the four different $L_{P}$. In the main text, we report the average test set MSE for each model across different $L_{P}$ on each dataset and provide the full results in §B.3.
@@
Adaptability to Longer $L_{I}$   Previous baseline models typically use $L_{I}$ between 96 and 720, as longer $L_{I}$ often leads to overfitting to long-term patterns, ignoring more important local effects (Zeng et al., 2023; Nie et al., 2022; Liu et al., 2024a). However, the next-step prediction and varying lookback inputs in pure AR/WAVE Transformers help the model focus on tokens closer to the next step, improving generalization. As shown in Table 3, increasing $L_{I}$ from 512 to 4096 improves pure AR/WAVE performance, demonstrating scalability and the ability to properly leverage long-term effects. Also, the WAVE structure consistently boosts AR model performance across different $L_{I}$.
@@
Visualization analysis   Fig. LABEL:fig:vis_test_loss shows test loss curves for different Pure AR/WAVE attention mechanisms on the Weather and ETTm1 datasets, with WAVE consistently outperforming AR in both convergence speed and final loss. Fig. 10 visualizes attention input sequence, AR weights, $\mathbf{B}$, and $\mathbf{\Theta}$ matrices of a test datapoint on Weather, showing how MA weights decouple local patterns, allowing AR weights to focus on cyclic and long-term patterns. Additional visualizations in Figs. LABEL:fig:vis_real_weights_additional_1LABEL:fig:vis_real_weights_additional_4 reinforce that there are important long-term stable seasonal patterns for AR weights to capture that should not be disrupted by applying forget gates or EMA. This explains why gated linear attention underperforms linear attention in our experiments.
@@ Appendix A Appendix
Appendix B Related works
@@
TSF Structures   The use of neural network structures for TSF has been widely explored (Hochreiter & Schmidhuber, 1997; Rangapuram et al., 2018; Salinas et al., 2020). Recently, many Transformer-based TSF models with encoder-only and encoder-decoder structures have emerged (Li et al., 2019a; Zhou et al., 2021; Wu et al., 2021; Zhang & Yan, 2023; Nie et al., 2022; Liu et al., 2024a). However, these complex Transformer architectures have not significantly outperformed simpler MLP or linear models (Zeng et al., 2023; Das et al., 2023; Xu et al., 2024; Lu et al., 2024). Additionally, these models struggle to handle short-term effects properly with longer lookback windows, where, paradoxically, longer inputs often lead to worse performance.
B.1 Datasets

### v4

@@ 1 Introduction
In recent years, autoregressive (AR) decoder-only Transformer-based models (Vaswani, 2017; Radford, 2018) have been widely used in sequence modeling tasks across fields such as NLP (Brown et al., 2020; Touvron et al., 2023), CV (Chen et al., 2020; Esser et al., 2021; Chang et al., 2022; Liu et al., 2024a), and audio (Borsos et al., 2023). This structure is well-suited for various sequential generation and prediction tasks. However, in typical sequence modeling tasks like time series forecasting (TSF), there has been less exploration of this architecture compared to other structures. Most of the best-performing recent TSF models are encoder-only Transformers (Liu et al., 2024b; Nie et al., 2022), MLPs (Das et al., 2023; Lu et al., 2024), or even linear models (Zeng et al., 2023; Xu et al., 2024). The few relevant discussions mainly focus on using pretrained autoregressive LLMs or similar structures for few-shot and zero-shot prediction (Gruver et al., 2023; Jin et al., 2024; Das et al., 2024; Liu et al., 2024c), with little research directly evaluating their TSF performance in end-to-end training. Therefore, this paper will first briefly demonstrate that with appropriate tokenization and training methods, a basic AR Transformer is enough to achieve results comparable to the state-of-the-art (SOTA) baselines, as shown in Fig. 1.
@@ 2.2 Appropriate tokenization for autoregressive forecasting
Recently, most time series forecasting research utilizes encoder-decoder or encoder-only Transformers for TSF (Li et al., 2019b; Zhou et al., 2021; Wu et al., 2021; Nie et al., 2022; Liu et al., 2024b), with limited focus on end-to-end decoder-only autoregressive Transformer because of error accumulation issue. For long-term forecasts, The autoregressive Transformers requires iteratively doing one-step prediction, leading to error accumulation and higher MSE compared to non-autoregressive models that generate the entire forecast at once.
@@ 2.5 Inspiration: decoupling the short-term impact
(a) Visualization of different effects with exponential decay strategies and their challenges in gated linear attention. (Left: a) Pure exponential decay strategy in gated linear attention; (Mid: b) Exponential decay facing challenges in capturing long-term dependencies; (Right: c) Exponential decay facing challenges in capturing periodic dependencies
However, applying exponential decay to the gated linear attention AR weights may not align with the needs of TSF, which often involves stable periodic patterns alongside short-term impacts. TSF data frequently exhibit seasonal effects that differ from the transient long-term effects in NLP. These seasonal effects are stable and persist across the temporal dimension without decaying. As shown in Figure 4(a), in scenarios involving seasonal effects, sudden changes, and local effects, relying solely on exponential decay in gated linear attention is not the most suitable approach for modeling TSF. Decoupling local effects into short-term MA weights, allowing the AR term to focus on modeling the seasonal and long-term effects it handles best, would be a better solution for TSF.
@@ 2.7 Indirect MA weight generation
With $\mathbf{\Theta}=\mathbf{B}\cdot(\mathbf{I}-\mathbf{B})^{-1}$, as long as the indirectly generated $\mathbf{\Theta}$ accurately reflects the characteristics of the MA weights we want, we can use $\sum_{j=1}^{t-1}\bm{\beta}_{t-1,j}\odot\bm{r}_{j}$ as $\bm{o}^{\text{MA}}_{t}$. Since $\bm{r}_{j}$ is known after computing $\bm{o}^{\text{AR}}_{t}$, linear attention can be used to compute $\bm{o}^{\text{MA}}_{t}$ without increasing the time complexity. To ensure the implicitly generated $\mathbf{\Theta}$ from $\mathbf{B}$ captures the desired MA properties, we must carefully design how $\mathbf{B}$ is generated. The invertibility of $(\mathbf{I}-\mathbf{B})^{-1}$ is guaranteed since $\mathbf{B}$ is strictly lower triangular. To efficiently compute the generated weights, we use the $\bm{\beta}_{t-1,j}=\phi_{q}^{\text{MA}}(\bm{q}^{\text{MA}}_{t-1})\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})$ to generate $\mathbf{B}$, similar to linear attention. Previous dynamic ARMA models in statistics often update MA weights based on observations (Grenier, 1983; Azrak and Mélard, 2006), so we derive $\bm{q}^{\text{MA}}_{t-1}$ and $\bm{k}^{\text{MA}}_{j}$ by multiplying the attention input $\bm{x}_{[\cdot]}$ with $\mathbf{W}^{\text{MA}}_{q}$ and $\mathbf{W}^{\text{MA}}_{k}$. Now, the effectiveness of MA weights lies in selecting the most suitable functions $\phi_{q}^{\text{MA}}(\cdot)$ and $\phi_{k}^{\text{MA}}(\cdot)$.
@@ 2.8 Selection of $\phi(\cdot)$ and characteristics of implicit MA weights
The MA term models short-term effects and local temporal relationships, so we want the implicit $\mathbf{\Theta}$ to follow a pattern where elements near the diagonal have larger absolute values, and those farther away gradually decrease. The expanded form of $\mathbf{\Theta}$ is given by $\mathbf{\Theta}=\mathbf{B}\cdot(\mathbf{I}-\mathbf{B})^{-1}=\mathbf{B}+\mathbf{B}^{2}+\mathbf{B}^{3}+\cdots$. The elements along the diagonal direction in $\mathbf{B}$ continually accumulate as products into the elements below them in $\mathbf{\Theta}$. Since $\mathbf{B}$ is strictly lower triangular, the elements of the subdiagonal in $\mathbf{\Theta}$ remain constant, while the elements further down progressively accumulate additional terms formed by the product of different $\beta_{[\cdot]}$ elements above. Assuming $\beta_{[\cdot]}$ follows a distribution and simplifying by setting each $\beta_{[\cdot]}$ to the distribution mean $b$, the elements of $\mathbf{\Theta}$ can be expressed as:
@@
We tested various activation function combinations for $\phi_{q}^{\text{MA}}(\cdot)$ and $\phi_{k}^{\text{MA}}(\cdot)$ to generate $\bm{\beta}_{t-1,j}=\phi_{q}^{\text{MA}}(\bm{q}^{\text{MA}}_{t-1})\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})$ values, as shown in Fig. 8(a). We used the sigmoid function $\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})=\sigma(\alpha\bm{k}^{\text{MA}}_{j}/\sqrt{d})$ to obtain values between 0 and 1, where $\alpha=0.05$  [^2] In the key activation, $\alpha$ controls the variance of each row in the $\mathbf{B}$ matrix, indirectly influencing the amount of long-term information (lower left) in the MA weights $\mathbf{\Theta}$. Increasing $\alpha$ would make the MA weights focus more on modeling long-term information. However, since we want the AR weights to handle the long-term component, we set $\alpha$ to a relatively small value. This explains why the rows of the $\mathbf{B}$ matrix appear smooth in the visualization. Refer to Fig. 22(a) for more details on $\alpha$, and see Fig. 22(b) for the effects of reversed positive $\phi_{q}$. and $\sqrt{d}$ are scaling factors to maintain small absolute values. Then, we selected a function $\phi_{q}^{\text{MA}}(\cdot)$ to make the product negative. We ultimately chose $\phi^{\text{MA}}_{q}(\bm{q}^{\text{MA}}_{t})=-\text{LeakyReLU}(-\bm{q}^{\text{MA}}_{t}/\sqrt{d})$ with a negative slope of 0.02. The inner negative sign maintains directional consistency (for later parameter sharing), and the outer negative sign encourages a negative output.
(a) Visualization of the $\mathbf{B}(\textrm{left})-\mathbf{\Theta}(\textrm{right})$ relationship with different $\phi(\cdot)$. We construct the simulated $\mathbf{B}$ matrices using randomly sampled $\bm{q}$ and $\bm{k}$ ($N=64$, $d=32$) from the normal distribution, and display the corresponding implicit $\mathbf{\Theta}$ matrices.
@@
Fig. 8(a) shows that LeakyReLU provides a balanced lag weight pattern. Unlike ReLU and Sigmoid, which only output values of the same sign, LeakyReLU offers some relaxation while keeping most values negative. This adds flexibility by enabling the desired negative smoothing effect of the MA term, with occasional positive values to enhance modeling flexibility.
@@ 3 Experiments
We conducted comprehensive experiments on 12 widely-used TSF datasets, including Weather, Solar, Electricity (ECL), ETTs, Traffic, and PEMS. See §A.2 for detailed description of datasets.
@@
Baselines We built AR Transformers using the five attention mechanisms from Table 1 and added MA terms to create WAVE attention for comparison in TSF tasks. Additionally, we included five recent SOTA baselines: FITS (Xu et al., 2024), iTransformer (Liu et al., 2024b), CATS (Lu et al., 2024), PatchTST (Nie et al., 2022), and DLinear (Zeng et al., 2023). We also used a simple channel-dependent encoder-only Transformer, modified by repeating the last input value (like NLinear) to address distribution shift. This model already surpasses older architectures like Autoformer (Wu et al., 2021) and Informer (Zhou et al., 2021), so we excluded these from our comparison.
In the main experiments, both pure AR and WAVE Transformers use a consistent setup: $m=3$ Transformer layers, 8 heads, and model dimension determined by a empirical method $d=16\sqrt{C}$, where $C$ is the number of series. We evaluate their performance using one-step prediction for each test datapoint, aligned with the baselines. Baseline hyperparameters are set to the reported values from their original papers. For more details on hyperparameters and implementation, see §A.3.
We ran all models on all datasets for the four different $L_{P}$. In the main text, we report the average test set MSE for each model across different $L_{P}$ on each dataset and provide the full results in §A.5.
@@
Adaptability to Longer $L_{I}$ Previous baseline models typically use $L_{I}$ between 96 and 720, as longer $L_{I}$ often leads to overfitting to long-term patterns, ignoring more important local effects (Zeng et al., 2023; Nie et al., 2022; Liu et al., 2024b). However, the next-step prediction and varying lookback inputs in pure AR/WAVE Transformers help the model focus on tokens closer to the next step, improving generalization. As shown in Table 3, increasing $L_{I}$ from 512 to 4096 improves pure AR/WAVE performance, demonstrating scalability and the ability to properly leverage long-term effects. Also, the WAVE structure consistently boosts AR model performance across different $L_{I}$.
@@
(a) Visualization of test loss curves. We show the testing performance of five attention mechanisms using pure AR/WAVE structures on the Weather and ETTm1 datasets ($L_{I}=512$, $L_{P}=48$).
@@
Visualization analysis Fig. 10(a) shows test loss curves for different Pure AR/WAVE attention mechanisms on the Weather and ETTm1 datasets, with WAVE consistently outperforming AR in both convergence speed and final loss. Fig. 10 visualizes attention input sequence, AR weights, $\mathbf{B}$, and $\mathbf{\Theta}$ matrices of a test datapoint on Weather, showing how MA weights decouple local patterns, allowing AR weights to focus on cyclic and long-term patterns. Additional visualizations in Figs. 22(c)22(f) reinforce that there are important long-term stable seasonal patterns for AR weights to capture that should not be disrupted by applying forget gates or EMA. This explains why gated linear attention underperforms linear attention in our experiments.
@@ Appendix A Appendix
A.1 Related works
Linear Attention Mechanisms The quadratic complexity of traditional attention has motivated extensive research into efficient alternatives. Katharopoulos et al. (2020) pioneered linear attention by replacing the exponential similarity function with kernel functions, achieving linear complexity through reordered matrix operations. While this enabled Transformers to be reformulated as RNNs during inference, early implementations suffered performance degradation. Various improvements followed: Choromanski et al. (2021) introduced Performers with FAVOR+ for unbiased softmax approximation; Qin et al. (2022) addressed unbounded gradients and attention dilution in TransNormer; and Sun et al. (2023) combined linear attention with retention mechanisms in RetNet. The connection to classical concepts was explored by Mao (2022), who linked linear transformers to Fast Weight Programmers from the 1990s. Recent advances focus on practical large-scale applications. Yang et al. (2024) proposed Gated Linear Attention with hardware-efficient training, while Lightning Attention-2 achieved constant training speed regardless of sequence length through innovative tiling strategies.
@@ A.1 Related works
Time Series Analysis Time series analysis encompasses several key tasks. Beyond forecasting, tasks include anomaly detection, which involves identifying abnormal points or patterns in data sequences (Chandola et al., 2009; Malhotra et al., 2015; Yang et al., 2025); classification, which assigns time series data to predefined categories or labels (Ismail Fawaz et al., 2019); clustering, grouping similar time series without predefined labels (Liao, 2005; Aghabozorgi et al., 2015); imputation, addressing missing data points to ensure continuity (Che et al., 2018); and change-point detection, pinpointing moments of significant shifts in statistical properties (Truong et al., 2020). Each of these tasks poses distinct challenges and requires tailored methodological approaches.
Time Series Forecasting Time series forecasting has evolved from classical methods like ARIMA (Box et al., 1974) and exponential smoothing (Holt, 2004) to deep learning approaches. RNN-based methods (Hochreiter and Schmidhuber, 1997; Rangapuram et al., 2018; Salinas et al., 2020) captured sequential dependencies but struggled with long-range patterns.
TSF Structures The use of neural network structures for TSF has been widely explored (Hochreiter and Schmidhuber, 1997; Rangapuram et al., 2018; Salinas et al., 2020; Wu et al., 2022). Recently, many Transformer-based TSF models with encoder-only and encoder-decoder structures have emerged (Li et al., 2019a; Zhou et al., 2021; Wu et al., 2021; Zhang and Yan, 2023; Nie et al., 2022; Liu et al., 2024b). Transformers revolutionized the field through various adaptations: LogTrans (Li et al., 2019b) with local convolutions, Informer (Zhou et al., 2021) with ProbSparse attention, and Autoformer (Wu et al., 2021) with auto-correlation mechanisms. However, these complex Transformer architectures have not significantly outperformed simpler MLP or linear models (Zeng et al., 2023; Das et al., 2023; Xu et al., 2024; Lu et al., 2024). Additionally, these models struggle to handle short-term effects properly with longer lookback windows, where, paradoxically, longer inputs often lead to worse performance. Recent work explores novel perspectives. Nie et al. (2022) proposed PatchTST treating time series as patches, while Liu et al. (2024b) applied attention across variates rather than time. The emergence of LLMs has opened new directions, with Gruver et al. (2023) demonstrating zero-shot forecasting capabilities and Jin et al. (2024) adapting pre-trained models to temporal tasks. Key challenges remain in balancing model complexity with performance and effectively modeling both temporal and cross-variate dependencies (Zhang and Yan, 2023; Lu et al., 2024), while integrating domain-specific biases without sacrificing generality.
A.2 Datasets